OpenStudy (anonymous):

Suppose a function f(x) has domain (-1,1). Find the domains of the following functions: A. f(x+1) B. f(1/x) C. f(squareroot x) D. f(x+1)/(x-1) Thanks !

OpenStudy (anonymous):

\[c) \sqrt{x}\]

OpenStudy (amistre64):

an odd question, but i think it has to do with transformations

OpenStudy (amistre64):

srry, i got no good ideas for this one :/

OpenStudy (anonymous):

ahh thanks anyways ; ]

OpenStudy (amistre64):

http://www.freemathhelp.com/forum/threads/87785-Domain-problem this seems to have a cogent solution process.

OpenStudy (amistre64):

if its good, then it seems like we take some f(a) such that: -1 < a < 1 and evaluate it ... f(1/x) would be like: -1 < 1/x < 1 x cant be a fraction, and it cant be 0, so x>1 or x < -1 seems to fit

OpenStudy (amistre64):

fine, it can be a fraction, but it has to be improper :) 1/(5/3) = 3/5 is fine 1/(2/3) = 3/2 is outside of the domain